Information from the financial statements of Ames Fabricators, Inc., included the following:
  

December 31
2018 2017
Common shares 100,000 100,000
Convertible preferred shares
(convertible into 38,000 shares of common)
15,600 15,600
10% convertible bonds
(convertible into 20,000 shares of common)
$ 1,000,000 $ 1,000,000


Ames’s net income for the year ended December 31, 2018, is $560,000. The income tax rate is 40%. Ames paid dividends of $5 per share on its preferred stock during 2018.

Required:

Compute basic and diluted earnings per share for the year ended December 31, 2018.
Answer #1

ANSWER

Basic earning per share = (net income-Preferred dividend)/Share outstanding = (560,000-15,600*5)/100,000 = 4.82 per share

Diluted earning per share = Adjusted net income/Adjusted shares

= 560,000+60,000/(100,000+38,000 +20,000)

=620,000/158,000

= 3.92 per share

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************